OpenStudy (anonymous):

Find the inverse f(x)=2x+4

OpenStudy (anonymous):

f(x) = 2x +4 For inverse, switch x & y x = 2*y +4 y = x/2 -2

OpenStudy (anonymous):

f(x) = 2x + 4 y = 2x + 4 y - 4 = 2x x = (y - 4)/2 g(y) = (y - 4)/2 g(x) = (x - 4)/2 where g(x) is the inverse